In 2017-18 the Pakistan Bureau of statistics announced that of all adult Pakistanis, 59,786,200 were employed, 3,613,800 were unemployed and 58,800,000 were not in the labor force. Use this information to calculate the labor-force participation rate.


1
Expert's answer
2021-05-05T14:23:34-0400

Employed Pakistan's =59786200


Unemployed Pakistan's=3613800


Not in the labor force=58800000


Total adult Pakistan's= Employed + Unemployed+ Not in the labor force


=59786200+3613800+58800000


=122200000


Labor force= Employed Pakistan's + Unemployed Pakistan's


=59786200+3613800


=63400000


"Labor force participation rate=\frac{labor force}{adult Pakistan's}× 100%"


"\frac{63400000}{122200000}×100%"


=51.8821%


=51.88%